### v2

@@ 2.4 Bayesian Instances
In this section, we show an improved analysis of the performance of Mechanism 1 under Bayesian settings. In particular, we consider a setting where for each agent $i\in\mathcal{N}$, his valuation function $v_{i}:2^{\mathcal{I}}\rightarrow\mathbb{R}_{\geq 0}$ is drawn from an underlying distribution $F_{i}$ satisfying the following assumption which we call the uniform highest value assumption: For each agent $i$, the probability that an item $j$ is agent $i$’s most valuable item is the same [^4] If there are multiple items that have the highest value, then the probability is counted only for one of the items. Our mechanism does not use this assumption and can be implemented regardless; we only use this assumption for the surplus guarantee for which it is sufficient that there exists a tie breaking among the items that satisfies the uniform highest value assumption. for all items, i.e., for all $j\in\mathcal{I}$ it holds that $\Pr[v_{i}(\{j\})=\max_{j^{\prime}\in\mathcal{I}}v_{i}(\{j^{\prime}\})]=\frac{1}{m}$. We also assume that the agents’ valuations are independent, and we denote by $\mathcal{F}=\mathcal{F}_{1}\times\dots\times\mathcal{F}_{n}$ the product of the agents’ distributions. In Bayesian settings, we are interested in maximizing the expected performance of a mechanism, and so we say that a mechanism $M$ guarantees an $\alpha$-approximation (for $\alpha\geq 1$) of the optimal welfare as residual surplus if:
@@
Our result in this section is that it is possible to extract a better fraction of the social welfare as residual surplus (Theorem 9), generalizing the results of [hartline2008optimal] from the case of identical items to heterogeneous items. Our result also captures the special case of unit-demand agents where each agent’s values are distributed identically and independently from the same distribution, which was studied in [goldner2024simple] concurrently and independently from our work.
@@
For a set $\mathcal{I}$ of $m$ heterogeneous items, $n$ unit-demand agents $\mathcal{N}$ with valuations sampled from the distribution $\mathcal{F}$ that satisfies the uniform highest value assumption, there exists a mechanism whose expected residual surplus is $O(\max\{1,\log\frac{n}{m}\})$ approximation to the welfare as residual surplus.
To prove this theorem, we first introduce some notation. Let $\mu_{v}:\mathcal{N}\to\mathcal{I}$ be a mapping from an agent to his most valuable item with respect to a valuation profile $v$, i.e., $\mu_{v}(i)=\arg\max_{j\in\mathcal{I}}v_{i}(\{j\})$. If an agent $i$ has more than one item with maximal value, we define $\mu_{v}(i)$ to be one of these items in a way that satisfies that $\Pr_{v_{i}\sim\mathcal{F}_{i}}[\mu_{v}(i)=j]=\frac{1}{m}$, for every item $j\in\mathcal{I}$. Moreover, we assume that agent $i$ breaks ties consistently in an independent way from $v_{-i}$. Similarly, let $\mu^{-1}_{v}:\mathcal{I}\to 2^{\mathcal{N}}$ be a mapping from an item to the set of agents whose most valuable item is this item with respect to a valuation profile $v$, i.e., $\mu^{-1}_{v}(j)=\{i\mid\mu_{v}(i)=j\}$.
@@
It holds that $\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]\geq\frac{\mbox{\bf E}{v\sim\mathcal{F}}[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})]}{\frac{e}{e-1}\cdot\max\{{1,\frac{n}{m}\}}}$.
@@
We prove this theorem by considering Mechanism 1 with parameters $r=\lceil\log_{2}(\frac{2e}{e-1}\cdot\max\{1,\frac{n}{m}\})\rceil,\,q=1$, and subroutine $B$ for unit demand agents (as defined in Unit-demand valuations.). Now, the expected residual surplus of this mechanism is:
@@
$\geq$ $\frac{1}{r+1}\left(\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]-\frac{(\frac{e}{e-1}\cdot\max\{1,\frac{n}{m}\})\cdot\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]}{2^{r}}\right)$
@@
where the first inequality is due to Lemma 3, the second inequality is since the social welfare of unit-demand agents is always bounded by giving each agent his favorite item, the third inequality is due to Lemma 10, and the last inequality is by assigning the value of $r$. ∎
@@
Consider the allocation rule $a$: For a valuation profile $v$, $a$ allocates each item $j$ to one of the agents in $\mu^{-1}_{v}(j)$ uniformly at random. To prove this lemma we bound the expected social welfare by the expected welfare of the allocation rule $a$. The expected welfare of an agent $i$ under allocation rule $a$ is equal to $\mbox{\bf E}_{v\sim\mathcal{F}}\left[\frac{v_{i}(\{\mu_{v}(i)\})}{|\mu^{-1}_{v}(\mu_{v}(i))|}\right]$. We next analyze the random variable of $|\mu^{-1}_{v}(\mu_{v}(i))|$. We know that by definition $i\in\mu^{-1}_{v}(\mu_{v}(i))$, and by the uniform highest value assumption we have that for every $i^{\prime}\in\mathcal{N}\setminus\{i\}$, it holds that $\Pr[i^{\prime}\in\mu^{-1}_{v}(\mu_{v}(i))]=\frac{1}{m}$. Combining with the independent assumption, we get that $\mid\mu^{-1}_{v}(\mu_{v}(i))\mid$ is distributed as $1+Bin(n-1,\frac{1}{m})$. Next we analyze the expected value of $\frac{1}{\mid\mu^{-1}_{v}(\mu_{v}(i))\mid}$.
$\mbox{\bf E}\left[\frac{1}{\mid\mu^{-1}_{v}(\mu_{v}(i))\mid}\right]&=\sum_{k=0}^{n-1}\frac{1}{k+1}\binom{n-1}{k}(\frac{1}{m})^{k}\cdot(1-\frac{1}{m})^{n-1-k}\\ &=\sum_{k=0}^{n-1}\frac{1}{n}\binom{n}{k+1}(\frac{1}{m})^{k}\cdot(1-\frac{1}{m})^{n-(k+1)}\\ &=\frac{m}{n}\sum_{k=0}^{n-1}\binom{n}{k+1}(\frac{1}{m})^{k+1}\cdot(1-\frac{1}{m})^{n-(k+1)}\\ &=\frac{m}{n}\cdot(1-(1-\frac{1}{m})^{n}),\end{aligned}$ (4)
where the second equality is due to the identity $\binom{n}{k}=\frac{n}{k}\binom{n-1}{k-1}$. The Last equality follows by observing that $\sum_{k=0}^{n-1}\binom{n}{k+1}(\frac{1}{m})^{k+1}\cdot(1-\frac{1}{m})^{n-(k+1)}$ is equal to the probability that a binomial random variable with parameters $n$ and $\frac{1}{m}$ equals to some $k$ between $1$ and $n$, i.e., it equals to the probability that a binomial random variable with parameters $n$ and $\frac{1}{m}$ is not equal to 00.
By the independence assumption we get that $v_{i}(\{\mu_{v}(i)\})$ is independent of $|\mu^{-1}_{v}(\mu_{v}(i))|$ which implies that:
$\mbox{\bf E}_{v\sim\mathcal{F}}\left[\frac{v_{i}(\{\mu_{v}(i)\})}{|\mu^{-1}_{v}(\mu_{v}(i))|}\right]=\mbox{\bf E}_{v\sim\mathcal{F}}[v_{i}(\{\mu_{v}(i)\})]\cdot\mbox{\bf E}_{v\sim\mathcal{F}}\left[\frac{1}{|\mu^{-1}_{v}(\mu_{v}(i))|}\right].$ (5)
Thus, by combining Equation 4 and Equation 5 we get that the expected welfare of the allocation rule $a$ is:
$\mbox{\bf E}_{v\sim\mathcal{F}}[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})]\cdot\frac{m}{n}\cdot(1-(1-\frac{1}{m})^{n}).$
@@
Case 1: $n\geq m$.
In this case $\frac{m}{n}\cdot(1-(1-\frac{1}{m})^{n})\geq\frac{m}{n}\cdot(1-(1-\frac{1}{m})^{m})\geq\frac{m}{n}\cdot(1-\frac{1}{e})$, and so we get:
$\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]\geq\mbox{\bf E}_{v\sim\mathcal{F}}[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})]\cdot\frac{m}{n}\cdot(1-\frac{1}{e}).$
Case 2: $n<m$.
In this case $\frac{m}{n}\cdot(1-(1-\frac{1}{m})^{n})\geq\frac{m}{n}\cdot(1-\frac{1}{e}^{\frac{n}{m}})$. Let $f(x)=x\cdot(1-e^{-\frac{1}{x}})$, and observe that this is an increasing function for $x\geq 0$. Since $n<m$, we have that $\frac{m}{n}\cdot(1-\frac{1}{e}^{\frac{n}{m}})>f(1)=1-\frac{1}{e}$. Overall:
$\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]\geq\mbox{\bf E}_{v\sim\mathcal{F}}[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})]\cdot\frac{m}{n}\cdot(1-(1-\frac{1}{m})^{n})\geq\mbox{\bf E}_{v\sim\mathcal{F}}[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})]\cdot(1-\frac{1}{e}).$
@@ 3 Divisible Goods
We now turn toward allocations with divisible goods. Since agents can be allocated fractions of each good, $\Delta$ is now the set of allocations which gives each agent $i$ a fraction $x_{ij}\in[0,1]$ of each good $j$ with the constraint that $\sum_{i\in\mathcal{N}}{x_{ij}}\leq 1,\forall j\in\mathcal{I}$. As before, each agent is associated with a valuation function $v_{i}:[0,1]^{|\mathcal{I}|}\rightarrow\mathbb{R}_{\geq 0}$ and we assume that valuation functions are monotone (i.e., $v_{i}(\vec{x})\leq v_{i}(\vec{y})$ for all $\vec{y}\geq\vec{x}$ where comparison between two vectors is coordinate-wise) and normalized (i.e., $v_{i}(\vec{0})=0$). In this section, we consider the case where agents valuation functions are concave. We say that a valuation function $v_{i}$ is concave if for any two allocation vectors $\vec{x}$ and $\vec{y}$ and any $\lambda\in[0,1]$ we have that $\lambda v_{i}(\vec{x})+(1-\lambda)v_{i}(\vec{y})\leq v_{i}(\lambda\vec{x}+(1-\lambda)\vec{y})$. Note that for any normalized and concave valuation function $v_{i}$ we have that $v_{i}(\lambda\vec{x})\geq\lambda v_{i}(\vec{x})$ for any $\lambda\in[0,1]$ and any allocation $\vec{x}$.
We adapt the VCG with copies framework introduced in Section 2.2 to the divisible items setting by proposing the restricted capacity VCG mechanism below. We then use this mechanism to devise a truthful mechanism with optimal guarantees for the divisible goods setting we study. Rather than randomly selecting a number of copies of each item, the restricted capacity VCG mechanism randomly chooses a “capacity” for each item, i.e., a maximum amount of each item that each agent may receive. The mechanism then runs VCG on a version of the input modified by this capacity constraint. This mechanism is, in a sense, simpler than Mechanism 1 since it circumvents the need for rounding any of the allocations output by VCG since they are all feasible on the initial instance (since items are divisible). As such, in this setting we are able to find a universally truthful mechanism, i.e., a mechanism which is a randomization over truthful mechanisms, which provides a stronger incentive guarantee than truthfulness in expectation.
Input: An items set $\mathcal{I}$, valuation profile $v_{1},\ldots,v_{n}$, parameters $r\in\mathbb{N}$
@@ 2 Let $q=2^{-\ell}$
Observe that in line 2 of Mechanism 2 one runs the VCG auction on a modified input valuations $\vec{v}^{q}$. This is equivalent to running a maximal-in-range mechanism on the original input valuations $\vec{v}$ but only over outcomes which allocate at most a $q$ fraction of any good to any agent. Maximal-in-range mechanisms are known to be truthful and ex-post individually rational (see, e.g., [nisan2007computationally]). But then, Mechanism 2 is a randomization over truthful and ex-post individually rational mechanisms and, hence, Mechanism 2 is universally truthful and ex-post individually rational. ∎
@@
It holds that $SW(\mathcal{N},q,\mathcal{I})-\frac{1}{2}SW(\mathcal{N},2q,\mathcal{I})\geq\frac{p(\mathcal{N},q,\mathcal{I})}{2}$.
@@
Mechanism 2 obtains expected residual surplus that is a $O\left(\log(n)\right)$ approximation to the social welfare.
@@
where the equality uses the fact that if allocation capacity of each item is capped at $1/n$ then the payments are 00 (since each agent has no externality) and the inequality applies Claim 14 to each term in the summation with $t\geq 1$. Finally, we may combine Equations (6) and (7) to obtain
@@
4 Improved Approximations for Two Agents

### v3

@@ 2.4 Bayesian Instances
[…]
@@
Our result in this section is that it is possible to extract a better fraction of the social welfare as residual surplus (Theorem 9). The case of $c=1$ of this result generalizes the case of identical items studied in [hartline2008optimal], and the case of unit-demand agents where each agent’s values are distributed identically and independently from the same distribution, which was studied in [goldner2024simple] independently from our work.
@@
For a set $\mathcal{I}$ of $m$ heterogeneous items, $n$ unit-demand agents $\mathcal{N}$ with valuations sampled from the distribution $\mathcal{F}$ that satisfies the no-superstar-item assumption for some constant $c\geq 1$, there exists a mechanism whose expected residual surplus is an $O(\log(1+\frac{n\cdot c}{m}))$ approximation to the welfare as residual surplus.
To prove this theorem, we first introduce some notation. Let $\mu_{v}:\mathcal{N}\to\mathcal{I}$ be a mapping from an agent to his most valuable item with respect to a valuation profile $v$, i.e., $\mu_{v}(i)=\arg\max_{j\in\mathcal{I}}v_{i}(\{j\})$. If an agent $i$ has more than one item with maximal value, we define $\mu_{v}(i)$ to be one of these items in a way that satisfies that $\Pr_{v_{i}\sim\mathcal{F}_{i}}[\mu_{v}(i)=j]\leq\frac{c}{m}$, for every item $j\in\mathcal{I}$. Moreover, we assume that agent $i$ breaks ties consistently in an independent way from $v_{-i}$. Similarly, let $\mu^{-1}_{v}:\mathcal{I}\to 2^{\mathcal{N}}$ be a mapping from an item to the set of agents whose most valuable item is this item with respect to a valuation profile $v$, i.e., $\mu^{-1}_{v}(j)=\{i\mid\mu_{v}(i)=j\}$.
@@
For a set $\mathcal{I}$ of $m$ heterogeneous items, $n$ unit-demand agents $\mathcal{N}$ with valuations sampled from the distribution $\mathcal{F}$ that satisfies the no-superstar-item assumption for some constant $c\geq 1$, it holds that
$\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]\geq\frac{\mbox{\bf E}_{v\sim\mathcal{F}}[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})]}{\frac{e}{e-1}\cdot\max\{{1,\frac{n\cdot c}{m}\}}}.$
@@
We prove this theorem by considering Mechanism 1 with parameters $r=\lceil\log_{2}(\frac{2e}{e-1}\cdot\max\{1,\frac{n\cdot c}{m}\})\rceil$, and rounding scheme $B_{UD}$ with $q=1$ (as defined in Unit-demand valuations.). Now, the expected residual surplus of this mechanism is:
@@
$\geq$ $\frac{1}{r+1}\left(\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]-\frac{(\frac{e}{e-1}\cdot\max\{1,\frac{n\cdot c}{m}\})\cdot\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]}{2^{r}}\right)$
@@
where the first inequality is due to Lemma 3, the second inequality is since the social welfare of unit-demand agents is always bounded by giving each agent his favorite item, the third inequality is due to Lemma 10, and the last inequality is by assigning the value of $r$.
The theorem holds since $r\in O(\log(1+\frac{n\cdot c}{m}))$. ∎
@@
Consider the allocation rule $a$: For a valuation profile $v$, $a$ allocates each item $j$ to one of the agents in $\mu^{-1}_{v}(j)$ uniformly at random. To prove this lemma we bound the expected social welfare by the expected welfare of the allocation rule $a$. The expected welfare of an agent $i$ under allocation rule $a$ is equal to $\mbox{\bf E}_{v\sim\mathcal{F}}\left[\frac{v_{i}(\{\mu_{v}(i)\})}{|\mu^{-1}_{v}(\mu_{v}(i))|}\right]$. We next analyze the random variable of $|\mu^{-1}_{v}(\mu_{v}(i))|$. We know that by definition $i\in\mu^{-1}_{v}(\mu_{v}(i))$, and by the no-superstar-item assumption we have that for every $i^{\prime}\in\mathcal{N}\setminus\{i\}$, it holds that $\Pr[i^{\prime}\in\mu^{-1}_{v}(\mu_{v}(i))]\leq\frac{c}{m}$. Combining with the independent assumption, we get that $|\mu^{-1}_{v}(\mu_{v}(i))|$ is $1+\sum_{i^{\prime}\neq i}\mathbb{I}[\mu_{v}(i^{\prime})=\mu_{v}(i)]$ which by our independence and no-superstar-item assumptions is stochastically dominated by $1+Bin(n-1,\frac{c}{m})$. Next we analyze the expected value of $\frac{1}{\mid\mu^{-1}_{v}(\mu_{v}(i))\mid}$.
$\mbox{\bf E}\left[\frac{1}{|\mu^{-1}_{v}(\mu_{v}(i))|}\right]$ $\geq\mbox{\bf E}\left[\frac{1}{1+Bin(n-1,\frac{c}{m})}\right]$
$=\sum_{k=0}^{n-1}\frac{1}{k+1}\binom{n-1}{k}\left(\frac{c}{m}\right)^{k}\cdot\left(1-\frac{c}{m}\right)^{n-1-k}$
$=\frac{m}{n\cdot c}\sum_{k=0}^{n-1}\binom{n}{k+1}\left(\frac{c}{m}\right)^{k+1}\cdot\left(1-\frac{c}{m}\right)^{n-(k+1)}$
$=\frac{m}{n\cdot c}\cdot\left(1-\left(1-\frac{c}{m}\right)^{n}\right),$ (5)
where the second equality is due to the identity $\binom{n}{k}=\frac{n}{k}\binom{n-1}{k-1}$. The last equality follows by observing that $\sum_{k=0}^{n-1}\binom{n}{k+1}(\frac{c}{m})^{k+1}\cdot(1-\frac{c}{m})^{n-(k+1)}$ is equal to the probability that a binomial random variable with parameters $n$ and $\frac{c}{m}$ equals to some $k$ between $1$ and $n$, i.e., it equals to the probability that a binomial random variable with parameters $n$ and $\frac{c}{m}$ is not equal to 00.
We note that since Inequality (5) holds for every realization of $\mu(i)$, it implies that:
$\mbox{\bf E}_{v\sim\mathcal{F}}\left[\frac{v_{i}(\{\mu_{v}(i)\})}{|\mu^{-1}_{v}(\mu_{v}(i))|}\right]\geq\mbox{\bf E}_{v\sim\mathcal{F}}[v_{i}(\{\mu_{v}(i)\})]\cdot\frac{m}{n\cdot c}\cdot\left(1-\left(1-\frac{c}{m}\right)^{n}\right).$
Thus, we get that the expected welfare of the allocation rule $a$ is:
$\mbox{\bf E}_{v\sim\mathcal{F}}\left[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})\right]\cdot\frac{m}{n\cdot c}\cdot\left(1-\left(1-\frac{c}{m}\right)^{n}\right).$
@@
Case 1: $n\cdot c\geq m$.
In this case $\frac{m}{n\cdot c}\cdot(1-(1-\frac{c}{m})^{n})\geq\frac{m}{n\cdot c}\cdot(1-(1-\frac{c}{m})^{\frac{m}{c}})\geq\frac{m}{n\cdot c}\cdot(1-\frac{1}{e})$, and so we get:
$\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]\geq\mbox{\bf E}_{v\sim\mathcal{F}}\left[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})\right]\cdot\frac{m}{n\cdot c}\cdot\left(1-\frac{1}{e}\right),$
which concludes the proof of this case.
Case 2: $n\cdot c<m$.
In this case $\frac{m}{n\cdot c}\cdot(1-(1-\frac{c}{m})^{n})\geq\frac{m}{n\cdot c}\cdot\left(1-{e}^{-\frac{cn}{m}}\right)$. Let $f(x)=x\cdot(1-e^{-\frac{1}{x}})$, and observe that this is an increasing function for $x>0$. Since $n\cdot c<m$, we have that
$\frac{m}{n\cdot c}\cdot\left(1-e^{\frac{-n\cdot c}{m}}\right)=f\left(\frac{m}{n\cdot c}\right)\geq f(1)=1-\frac{1}{e}.$
Overall for both cases:
$\mbox{\bf E}_{v\sim\mathcal{F}}[SW(\mathcal{N},\mathcal{I})]\geq\mbox{\bf E}_{v\sim\mathcal{F}}\left[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})\right]\cdot\frac{m}{n\cdot c}\cdot\left(1-\left(1-\frac{c}{m}\right)^{n}\right)\geq\frac{\mbox{\bf E}_{v\sim\mathcal{F}}\left[\sum_{i\in\mathcal{N}}v_{i}(\{\mu_{v}(i)\})\right]}{\frac{e}{e-1}\cdot\max\{1,\frac{n\cdot c}{m}\}}.$
@@ 3 Divisible Goods
We now turn toward allocations with divisible goods. Since agents can be allocated fractions of each good, $\Delta$ is now the set of allocations which gives each agent $i$ a fraction $x_{ij}\in[0,1]$ of each good $j$ with the constraint that $\sum_{i\in\mathcal{N}}{x_{ij}}\leq 1,\forall j\in\mathcal{I}$. As before, each agent is associated with a valuation function $v_{i}:[0,1]^{|\mathcal{I}|}\rightarrow\mathbb{R}_{\geq 0}$ and we assume that valuation functions are monotone (i.e., $v_{i}(\vec{x})\leq v_{i}(\vec{y})$ for all $\vec{y}\geq\vec{x}$ where comparison between two vectors is coordinate-wise) and normalized (i.e., $v_{i}(\vec{0})=0$). In this section, we consider the case where agents’ valuation functions are concave. We say that a valuation function $v_{i}$ is concave if for any two allocation vectors $\vec{x}$ and $\vec{y}$ and any $\lambda\in[0,1]$ we have that $\lambda v_{i}(\vec{x})+(1-\lambda)v_{i}(\vec{y})\leq v_{i}(\lambda\vec{x}+(1-\lambda)\vec{y})$. Note that for any normalized and concave valuation function $v_{i}$ we have that $v_{i}(\lambda\vec{x})\geq\lambda v_{i}(\vec{x})$ for any $\lambda\in[0,1]$ and any allocation $\vec{x}$.
We adapt the VCG with copies framework introduced in Section 2.2 to the divisible items setting by proposing the restricted capacity VCG mechanism below. We then use this mechanism to devise a truthful mechanism with optimal guarantees for the divisible goods setting we study. Rather than randomly selecting a number of copies of each item, the restricted capacity VCG mechanism randomly chooses a “capacity” $q$ for each item, i.e., a maximum amount $q$ of each item that each agent may receive. The mechanism then runs VCG on a version of the input modified by this capacity constraint. This mechanism is, in a sense, simpler than Mechanism 1 since it circumvents the need for rounding any of the allocations output by VCG since they are all feasible on the initial instance (since items are divisible). As such, in this setting we are able to find a universally truthful mechanism, i.e., a mechanism which is a randomization over truthful mechanisms, which provides a stronger incentive guarantee than truthfulness in expectation.
Input: An item set $\mathcal{I}$, valuation profile $v_{1},\ldots,v_{n}$, parameters $r\in\mathbb{N}$
@@ 2 Let $q=2^{-\ell}$
Observe that in line 2 of Mechanism 2 one runs the VCG auction on a modified input valuations $\vec{v}^{q}$. This is equivalent to running a maximal-in-range mechanism on the original input valuations $\vec{v}$ but only over outcomes that allocate at most a $q$ fraction of any good to any agent. Maximal-in-range mechanisms are known to be truthful and ex-post individually rational (see, e.g., [nisan2007computationally]). But then, Mechanism 2 is a randomization over truthful and ex-post individually rational mechanisms and, hence, Mechanism 2 is universally truthful and ex-post individually rational. ∎
@@
For agents with concave valuations it holds that $SW(\mathcal{N},q,\mathcal{I})-\frac{1}{2}SW(\mathcal{N},2q,\mathcal{I})\geq\frac{p(\mathcal{N},q,\mathcal{I})}{2}$.
@@
When Mechanism 2 receives concave valuations it obtains expected residual surplus that is a $O\left(\log(n)\right)$ approximation to the social welfare.
We again note that the lower bound of hartline2008optimal implies a lower bound of $\Omega(\log(n))$ for buyers with concave valuation functions (see Section 1.1), therefore the result of Theorem 15 asymptotically tight.
@@
where the equality uses the fact that if the allocation capacity of each item is capped at $1/n$ then the payments are 00 (since each agent has no externality) and the inequality applies Claim 14 to each term in the summation with $t\geq 1$. Finally, we may combine Inequalities (6) and (7) to obtain
@@
4 Discussion and Future Directions
In this paper, we study mechanism-design for maximizing consumer surplus. Our work is the first to achieve a sub-linear approximation to the welfare in adversarial multi-parameter settings, and we obtain asymptotically optimal guarantees for various settings including an $O(\log(n))$-approximation for unit-demand and multi-unit submodular agents, an $O(\log{(1+\frac{n\cdot c}{m})})$-approximation for Bayesian unit-demand settings satisfying the no-superstar-item assumption, and an $O(\log(n))$-approximation for agents with concave valuations over divisible goods. We also resolve some additional open questions regarding instances with two bidders from [hartline2008optimal].
We remark that all our mechanisms are prior-independent (or prior-free) while the matching lower bounds are achieved in instances where the agents’ values are independently sampled from the same distribution and the mechanisms are only required to be Bayesian incentive compatible. A natural future direction is whether this is true for other settings as well, whether Bayesian incentive compatible mechanisms have the same worst-case guarantees as prior-independent ones.
Beyond Unit-Demand and Multi-Unit.
We next discuss how to adapt Mechanism 1 for a broader class of valuations. Consider the case where agents’ valuations are weighted matroid valuation functions under known matroids. In other words, each agent $i\in\mathcal{N}$ is associated with a (publicly known to the auctioneer) matroid $M_{i}=(\mathcal{I},E_{i})$, and private values $v^{i}_{1},\ldots,v^{i}_{m}\in\mathbb{R}_{\geq 0}$ where agent’s $i$ valuation function $v_{i}:2^{\mathcal{I}}\rightarrow\mathbb{R}_{\geq 0}$ satisfies $v_{i}(S)=\max_{T\in E_{i}}\sum_{j\in S\cap T}v^{i}_{j}$. An allocation $A^{\prime}=(A^{\prime}_{1},\dots A^{\prime}_{n})$ is non-redundant with respect to the class of weighted matroid functions (for known matroids) if $g(A^{\prime}_{i})\in E_{i}$ and an agent is not allocated two copied of the same item. For this case, we construct a $1$-valid rounding scheme that allocates each item independently with probability $\frac{1}{2^{\ell}}$ to one of its recipients of the copies. This is indeed a $1$-valid rounding scheme since matroids are downward-closed, which implies that the expected value each agent receives by independent allocation satisfies Equation (1).
Mechanism  1 applied with this rounding scheme is TIE and ex-post IR by Claim 1, and provides an $O(\log n)$ approximation by Lemma 3 as weighted matroid functions are gross-substitutes, and since weighted matroid functions are closed [^7] By Claim 4 under the transformation that creates copies, and by similar arguments the same holds with respect to weighted matroid functions. under the operation of adding copies of items. For achieving the $O(\log n)$-approximation, the mechanism needs to know the underlying matroids $M_{i}$ (if the matroids are private information and the agents report them, then the truthfulness proof requires that the rounding scheme should be applied to all possible allocations and not just allocations that allocate feasible sets to the agents).
We leave as an open direction to extend our framework (or variants of it) beyond the settings studied in this paper, or in general devise truthful mechanisms with optimal surplus guarantees for other classes of functions (e.g., submodular, XOS, subadditive, supermodular, and additive subject to downward-closed constraints).
An alternative Benchmark
In this paper, we considered mechanism design for surplus maximization and compared it to the first-best benchmark, i.e., the social welfare. An interesting open question is establishing guarantees with respect to tighter benchmarks for surplus maximization. One option is trying to generalize the $\mathcal{G}$ benchmark introduced in [hartline2008optimal] for single-parameter settings. However, the definition of $\mathcal{G}$ relies heavily on the single-parameter structure, and it is not clear whether there is an intuitive way to generalize it to multi-dimensional settings. Furthermore, as discussed in [goldner2024simple], finding a tractable alternative benchmark seems to be a difficult problem in its own right. They demonstrate this by examining two prominent approaches in revenue maximization - bounding the optimal revenue via Langrangian duality [duality-paper] or by generating copies of the agents [Chawla-copies] - and highlighting the challenges in applying these methods to multi-dimensional surplus maximization settings.
Appendix A Omitted Proofs
A.1 Proof of Claim 14
Proof of Claim 14.
$\sum_{i\in\mathcal{N}}{p_{i}(\mathcal{N},q,\mathcal{I})}\leq\sum_{i\in\mathcal{N}}{SW(\mathcal{N},q,\mathcal{I}+A_{i}(\mathcal{N},q,\mathcal{I}))-SW(\mathcal{N},q,\mathcal{I})}.$ (8)
where the first inequality is due to the fact that the resulting allocation from giving each agent exactly half of what she receives in $SW(\mathcal{N},2q,\mathcal{I})$ is a feasible allocation when the item input is $\mathcal{I}/2$ and capacity is $q$, the first equality is by telescoping sum, the second inequality is due to the fact that by Claims 12 and 13 we have that the social welfare function is concave, and the final inequality is also due to this fact. Combining Inequalities (8) and (9) then completes the proof of the claim. ∎
Appendix B Improved Approximations for Two Agents